# Apollo: An Lightweight Multilingual Medical LLM towards Democratizing Medical AI to 6B People

### v4

@@ 1 Introduction
Rationale of Multilinguality in Medical LLMs The rationale of multilingual adaption in medical LLMs (Cox and Maryns, 2021; Pecina et al., 2014) could be twofold. Firstly, non-native English-speaking doctors often engage in bilingual learning through their native language and English from the outset, naturally introducing multilingual challenges in the learning process (Markó et al., 2006). Secondly, to better serve local communities, especially in countries and regions with scarce medical resources, medical aid based on local languages often achieves higher communication efficiency and acceptance (Brindley et al., 2014; Albrecht et al., 2013). Meanwhile, local medical knowledge can complement mainstream medical knowledge, fostering mutual benefits to accelerate medical development (Klayman, 1985; Yuan et al., 2016). Our pilot study in Sec. 2 also reveals that joint training of multiple languages enhances performance in the medical LLMs, indicating a beneficial complementarity among languages.
@@
The Corpora: ApolloCorpora Towards building multilingual medical LLMs, the first step is to build high-qaulity corpora. We select the six most populous languages: English, Chinese, Hindi, Spanish, French, and Arabic for experiments  [^1] See the language popularity in https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers, The set of the selected six languages covers a total of 6.1 billion people in 132 countries and regions, according to https://en.wikipedia.org/wiki/List_of_countries_and_dependencies_by_population.. As shown in Fig. 1, it encompasses a diverse array of linguistic backgrounds, particularly in regions often characterized by limited medical resources (e.g., areas with lower average life expectancies). We collect and process data from data sources including books, clinical guidelines, encyclopedias, papers, online forums and examinations, obtaining ApolloCorpora with 2.5B tokens.
@@
The Lite LLM: Apollo The resulted multilingual medical LLMs trained by ApolloCorpora are named Apollo, to commemorate the Greek deity associated with healing, disease, the Sun, and light; this symbolizes the democratization of medical LLMs to 6 billion people, illuminating global healthcare. We explore a new domain adaption method that rewrite the pre-training corpora into QA pairs using ChatGPT (Li et al., 2023c) and adaptively sample training data, resulting in a smoother transition compared with the typical paradigm with continued pretraining and instruction tuning. Apollo ranges from 2B to 7B parameters. The advantage of the relatively-small model scale includes potential use as draft models for speculative decoding (Leviathan et al., 2023) or as proxy models for proxy-tuning (Liu et al., 2024a). In particular, we apply proxy-tuning on top of Apollo to larger general LLMs, enhancing its multilingual medical capabilities. This is achieved without the need to directly train the general model using sensitive medical corpora, thereby underscoring the practical significance of Apollo in terms of protecting the privacy of medical training data against centralized training methods.
@@
The Benchmark: XMedBench We select local multiple-choice tasks to assess models’ medical knowledge. For Hindi and Arabic, which lack local assessments, we choose to translate the medical-related parts of MMLU (Hendrycks et al., 2020). The results show that the gap between open source and closed source is narrowing. While GPT-4 has demonstrated superior efficacy across numerous languages, the Apollo series models achieve the best performance among models of equivalent size.
@@
Contributions The contributions of this paper are as follows. (1) We collect and organize a high-quality multilingual medical corpora ApolloCorpora with rich local characteristics; (2) We obtain a series of SOTA multilingual LLMs Apollo at various parameter scales (especially in relatively-small sizes) ; (3) By using proxy-tuning, Apollo could significantly improve larger general LLMs without finetuning, providing a new way to mitigating the exposure of private medical training data to centralized training systems; (4) We introduce a multilingual medical evaluation XMedBench and conduct extensive benchmark of existing models.
@@
2 Pilot Study on the Multilinguality of Medical LLMs
2.1 The Research Question
Language-neutral Hypothesis It is commonly believed that knowledge, whether medical or general, should be independent of language. For example, the fact that the sun rises in the east remains unchanged whether expressed in English or Chinese, suggesting that knowledge might be considered language-neutral. Consequently, medical corpora in various languages could serve as an augmentation for training, thereby enhancing the efficacy of the resulting medical LLMs.
Language-dependent Hypothesis However, due to historical, cultural, and regional political influences, medical knowledge can vary significantly across different cultural contexts, especially as reflected in language. The integration of medical knowledge across languages might dilute the local specificity of medicine due to differences in lifestyle and constitution across regions (Rotti et al., 2014; Sharma et al., 2020). For instance, in traditional Chinese medicine, colds are classified into types caused by heat or cold, and treatments may vary locally, relying on unique medical recipe that uses local herbs or substances. This indicates that historical and cultural traditions can shape medical knowledge to some extent.
Can medical data in different languages complement or harm each?
2.2 The Pilot Study
Experimental settings To investigate the above question, we use a lite multilingual LLM Qwen-1.8B (Bai et al., 2023) as the LM backbone. It is argued that the findings should be agnostic to the selection of LM backbones, the selection of LLM Qwen-1.8B is due to its popularity, performance and more the importantly multilingual support. In the monolingual training setting, the LM backbone is further trained by individual language, resulted six language-specific LLM variants (i.e. , English, Chinese, French, Spanish, Arabic, and Hindi.). The training data used could be found in Sec. 3. Moreover, we average the weights of these six LLM variants and obtain a new model that does not need further training, denoted as ‘weight average’. In the multilingual training setting, we train the LM using a mixture of the corpora in these six languages.
@@ 2.2 The Pilot Study
Books Continue pre-training EN (296.7M), ZH (117.1M) 413.8M
Papers Continue pre-training ZH (45.6M), EN (252.9M), ES (46.0M), FR (4.5M) 349.0M
Encyclopedias Continue pre-training EN (221.1M), FR (4.6M), HI (0.5M) 226.2M
Dialogues Continue pre-training EN (92.1M), ZH (46.6M), AR (10.4M) 149.1M
@@
Guidelines Continue pre-training EN (29.6M) 29.6M
@@
Findings. Tab. 1 highlights the effectiveness of our methods in leveraging multilingual data to enhance the performance of medical Large Language Models (LLMs). The language-specific models, each trained exclusively on data from one language, demonstrate varying degrees of improvement in their respective languages over the original LM, underscoring the value of language-specific training. However, these models show limitations outside their target languages, as seen in the relatively low scores in non-target languages, particularly in the Apollo-French-1.8B and Apollo-Hindi-1.8B models. Our method, which includes both the weight average and mmultilingual training, significantly outperforms language-specific models across all languages in terms of average performance, as shown in the last column of Tab. 1. This illustrates the substantial benefits of combining multilingual data for training medical LLMs, with marked improvements in understanding and generating medical information across a diverse set of languages.
@@
Figure 2: Overview of this work, including corpora, a benchmark and a series of medical LLMs.
Figure 3: Prompt Template for XMedBench
@@ 3.1 ApolloCorpora: Data Collection and Cleaning
Data Leakage Checking The issue of data leakage is a recent focus of the academic community, which largely determines whether the results of the paper are convincing. For knowledge embedding tasks, data leakage screening with different stringency often leads to different performance. We follow Med-PaLM2 (Singhal et al., 2023) and adopt a more stringent deletion strategy. Specifically, we define a data item as leaked data if the entire question or at least 64 consecutive characters overlap with the data item. Regarding the exam exercise data source, there were 580,645 exercises before screening, and 3,041 exercises are deleted, with a screening rate of 0.52%. For other data sources, since they are not exam questions, there is no difference before and after filtering.
@@ 3.2 Apollo, the Lite Multilingual Medical LLM
Mix Training Our dataset $D$ comprises continuing pre-training data $D_{PT}$ and instruction tuning data $D_{SFT}$. The sampling probability of each data $x\in D$ changes during training. The sampling probability of data $x$ at step $t$ during training was determined using priority sampling, defined as:
@@
Settings We set the priority $\pi(x)=16$ for $x\in D_{PT}$, and $\pi(x)=2$ for $x\in D_{SFT}$. In order to achieve the purpose of smooth transition of sampling ratio. The overall training sequence of pre-training corpus first, and then instruction Tuning corpus is maintained, but the transition can be smoothed. The Batch size of model training is set to 256, the learning rate is set to 1e-5, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the instruction data is trained for two epochs.
@@ 4.1 XMedBench: Multilingual Medical Knowledge Evaluation
Construction of XMedBench For English, we use the MedQA-USMLE (Zhang et al., 2018), MedMCQA (Pal et al., 2022), and medical-related parts of MMLU (Hendrycks et al., 2020); for Chinese, we used the of MedQA-MCMLE (Zhang et al., 2018) and medical-related parts CMMLU (Li et al., 2023a); for Spanish, we used HEAD-QA (Vilares and Gómez-Rodríguez, 2019); for French, we used FrenMedMCQA (Labrak et al., 2023a); For Arabic and Hindi, which lack local evaluation, we compromised and followed Llama3’s multilingual evaluation method (et al., 2024), using Google Translate and inviting practicing physicians to proofread and finally get the translated version of MMLU. Specifically, we follow Med-PaLM2 (Singhal et al., 2023) and select six subcategories in MMLU: Clinical knowledge, Medical genetics, Anatomy, Professional medicine, College biology, and College medicine. For MedQA, we choose the 4-options version. For CMMLU, we select seven subdirectories: Anatomy, Clinical knowledge, College medicine, Genetics, Nutrition, Traditional chinese medicine, and Virology.
Settings We adopt 3-shot evaluation and use regular matching to extract options. The specific evaluation prompts are shown in Fig. 3. For the generation strategy, we do not perform sampling and set the maximum and minimum number of generated tokens to 128 and 2. For model loading, except for the 0.5B size model which uses full precision loading, we uniformly use half precision loading. Please see the App. B.2 for details of Models.
@@ 5 The Application of the Lite Apollo: Proxy-Tuning for Larger Models
Preliminaries Inspired by  Liu et al. (2024a, 2021), we introduce a lightweight model-agnostic decoding method in medical senarios. We leverage the logits from both pre and post fine-tuned small models to indirectly steer the larger base model’s adjustments, thereby eschewing the need for direct parameter fine-tuning. Let $M_{raw}$ denote the smaller pre-trained model, and $M_{tuned}$ denote its fine-tuned counterpart. We compute the logit offset as "proxy" for each token, corresponding to the anti-expert and expert roles as delineated in  Liu et al. (2021). This offset is then applied to the base model $M_{base}$ to synchronize the predictive distributions of the smaller and larger models. The modified probability distribution is given by:
@@
Settings $M_{base}$ is designated as the subject of investigation for Qwen-7B. Apollo-1.8B and Qwen-1.8B are appointed as the $M_{tuned}$ and $M_{raw}$ respectively.
Results As shown in the Tab. 5, the overall effect of the model improves a lot without changing the parameters after proxy-tuning. From language perspective, except English, all other languages increase, and French has the most obvious increase. Excitingly, for French and Spanish, the model after proxy-tuning performs better than both $M_{tuned}$ and $M_{base}$, indicating that new accurate knowledge is generated after proxy-tuning. We also notice a decline in English proficiency. This may be because there is a gap between the distribution of difference and the probability itself, which leads to over-strengthening of the second option and requires further exploration and optimization.
@@ A.3 Localized features of ApolloCorpora
Language English Chinese French Spanish Arabic Hindi Avg.
[…]
Our Models and their Bases
Qwen-0.5B 24.43 3.78 16.94 14.16 10.88 23.68 26.02 26.29 26.35 19.17
Qwen-1.8B 26.79 31.05 40.89 44.28 35.86 22.12 27.43 23.71 8.82 28.99
Gemma-2B 30.24 32.27 37.35 25.98 28.06 25.86 32.43 20.96 25.53 28.74
Apollo-2B 38.33 42.00 52.89 46.76 36.76 38.32 41.28 31.62 31.50 39.94
Yi-6B 45.48 47.98 62.27 78.90 69.47 45.79 47.01 12.22 10.74 46.65
Apollo-6B 56.25 57.53 68.65 85.52 72.62 51.71 58.47 33.46 33.61 57.54
Gemma-7B 53.42 50.94 70.15 48.95 43.29 57.63 62.79 36.21 48.58 52.44
Apollo-7B 56.00 58.21 71.86 72.36 59.04 60.44 63.73 41.82 45.55 58.78
Table 6: Model performance comparison before and after Mix Training

### v5

@@ 1 Introduction
Figure 2: Overview of this work, including corpora, benchmark, models and their application.
Rationale of Multilinguality in Medical LLMs
@@
The rationale of multilingual adaption in medical LLMs (Cox and Maryns, 2021; Pecina et al., 2014) could be twofold. Firstly, non-native English-speaking doctors often engage in bilingual learning through their native language and English from the outset, naturally introducing multilingual challenges in the learning process (Markó et al., 2006). Secondly, to better serve local communities, especially in countries and regions with scarce medical resources, medical aid based on local languages often achieves higher communication efficiency and acceptance (Brindley et al., 2014; Albrecht et al., 2013). Meanwhile, local medical knowledge can complement mainstream medical knowledge, fostering mutual benefits to accelerate medical development (Klayman, 1985; Yuan et al., 2016). Our pilot study in Sec. 2 also reveals that joint training of multiple languages enhances performance in the medical LLMs, indicating a beneficial complementarity among languages.
The Corpora: ApolloCorpora
@@
Towards building multilingual medical LLMs, the first step is to build high-qaulity corpora. We select the six most populous languages: English, Chinese, Hindi, Spanish, French, and Arabic for experiments  [^1] See the language popularity in https://en.wikipedia.org/wiki/List_of_languages_by_total_number_of_speakers, The set of the selected six languages covers a total of 6.1 billion people in 132 countries and regions, according to https://en.wikipedia.org/wiki/List_of_countries_and_dependencies_by_population.. As shown in Fig. 1, it encompasses a diverse array of linguistic backgrounds, particularly in regions often characterized by limited medical resources (e.g., areas with lower average life expectancies). We collect and process data from data sources including books, clinical guidelines, encyclopedias, papers, online forums and examinations, obtaining ApolloCorpora with 2.5B tokens.
The Lite LLM: Apollo
@@
The resulted multilingual medical LLMs trained by ApolloCorpora are named Apollo, to commemorate the Greek deity associated with healing, disease, the Sun, and light; this symbolizes the democratization of medical LLMs to 6 billion people, illuminating global healthcare. We explore a new domain adaption method that rewrite the pre-training corpora into QA pairs using ChatGPT (Li et al., 2023c) and adaptively sample training data, resulting in a smoother transition compared with the typical paradigm with continued pretraining and instruction tuning. Apollo ranges from 2B to 7B parameters. The advantage of the relatively-small model scale includes potential use as draft models for speculative decoding (Leviathan et al., 2023) or as proxy models for proxy-tuning (Liu et al., 2024a). In particular, we apply proxy-tuning on top of Apollo to larger general LLMs, enhancing its multilingual medical capabilities. This is achieved without the need to directly train the general model using sensitive medical corpora, thereby underscoring the practical significance of Apollo in terms of protecting the privacy of medical training data against centralized training methods.
The Benchmark: XMedBench
@@
We select local multiple-choice tasks to assess models’ medical knowledge. For Hindi and Arabic, which lack local assessments, we choose to translate the medical-related parts of MMLU (Hendrycks et al., 2020). The results show that the gap between open source and closed source is narrowing. While GPT-4 has demonstrated superior efficacy across numerous languages, the Apollo series models achieve the best performance among models of equivalent size.
Contributions
@@
As shown in Fig. 2, the contributions of this paper are as follows. (1) We collect and organize a high-quality multilingual medical corpora ApolloCorpora with rich local characteristics; (2) We obtain a series of SOTA multilingual LLMs Apollo at various parameter scales (especially in relatively-small sizes) ; (3) By using proxy-tuning, Apollo could significantly improve larger general LLMs without finetuning, providing a new way to mitigating the exposure of private medical training data to centralized training systems; (4) We introduce a multilingual medical evaluation XMedBench and conduct extensive benchmark of existing models.
2 Pilot Study on the Multilinguality of Medical LLMs
2.1 The Research Question
Language-neutral Hypothesis
It is commonly believed that knowledge, whether medical or general, should be independent of language. For example, the fact that the sun rises in the east remains unchanged whether expressed in English or Chinese, suggesting that knowledge might be considered language-neutral. Consequently, medical corpora in various languages could serve as an augmentation for training, thereby enhancing the efficacy of the resulting medical LLMs.
Language-dependent Hypothesis
However, due to historical, cultural, and regional political influences, medical knowledge can vary significantly across different cultural contexts, especially as reflected in language. The integration of medical knowledge across languages might dilute the local specificity of medicine due to differences in lifestyle and constitution across regions (Rotti et al., 2014; Sharma et al., 2020). For instance, in traditional Chinese medicine, colds are classified into types caused by heat or cold, and treatments may vary locally, relying on unique medical recipe that uses local herbs or substances. This indicates that historical and cultural traditions can shape medical knowledge to some extent.
Can medical data in different languages complement or harm each?
2.2 The Pilot Study
Experimental settings
To investigate the above question, we use a lite multilingual LLM Qwen-1.8B (Bai et al., 2023) as the LM backbone. It is argued that the findings should be agnostic to the selection of LM backbones, the selection of LLM Qwen-1.8B is due to its popularity, performance and more the importantly multilingual support. In the monolingual training setting, the LM backbone is further trained by individual language, resulted six language-specific LLM variants (i.e. , English, Chinese, French, Spanish, Arabic, and Hindi.). The training data used could be found in Sec. 3. Moreover, we average the weights of these six LLM variants and obtain a new model that does not need further training, denoted as ‘weight average’. In the multilingual training setting, we train the LM using a mixture of the corpora in these six languages.
@@ 2.2 The Pilot Study
Books Continue Pretrain EN (296.7M), ZH (117.1M) 413.8M
Papers Continue Pretrain ZH (45.6M), EN (252.9M), ES (46.0M), FR (4.5M) 349.0M
Encyclopedias Continue Pre-train EN (221.1M), FR (4.6M), HI (0.5M) 226.2M
Dialogues Continue Pretrain EN (92.1M), ZH (46.6M), AR (10.4M) 149.1M
@@
Guidelines Continue Pretrain EN (29.6M) 29.6M
@@
Findings
Tab. 1 highlights the effectiveness of our methods in leveraging multilingual data to enhance the performance of medical Large Language Models (LLMs). The language-specific models, each trained exclusively on data from one language, demonstrate varying degrees of improvement in their respective languages over the original LM, underscoring the value of language-specific training. However, these models show limitations outside their target languages, as seen in the relatively low scores in non-target languages, particularly in the Apollo-French-1.8B and Apollo-Hindi-1.8B models. Our method, which includes both the weight average and mmultilingual training, significantly outperforms language-specific models across all languages in terms of average performance, as shown in the last column of Tab. 1. This illustrates the substantial benefits of combining multilingual data for training medical LLMs, with marked improvements in understanding and generating medical information across a diverse set of languages.
@@ 3.1 ApolloCorpora: Data Collection and Cleaning
Data Leakage Checking
The issue of data leakage is a recent focus of the academic community, which largely determines whether the results of the paper are convincing. For knowledge embedding tasks, data leakage screening with different stringency often leads to different performance. We follow Med-PaLM2 (Singhal et al., 2023) and adopt a more stringent deletion strategy. Specifically, we define a data item as leaked data if the entire question or at least 64 consecutive characters overlap with the data item. Regarding the exam exercise data source, there were 580,645 exercises before screening, and 3,041 exercises are deleted, with a screening rate of 0.52%. For other data sources, since they are not exam questions, there is no difference before and after filtering.
@@ 3.2 Apollo, the Lite Multilingual Medical LLM
Mix Training
Our dataset $D$ comprises continuing pre-training data $D_{PT}$ and instruction tuning data $D_{SFT}$. The sampling probability of each data $x\in D$ changes during training. The sampling probability of data $x$ at step $t$ during training was determined using priority sampling, defined as:
@@
Settings
We set the priority $\pi(x)=16$ for $x\in D_{PT}$, and $\pi(x)=2$ for $x\in D_{SFT}$. In order to achieve the purpose of smooth transition of sampling ratio. The overall training sequence of pre-training corpus first, and then instruction Tuning corpus is maintained, but the transition can be smoothed. The Batch size of model training is set to 256, the learning rate is set to 1e-5, and the warm up rate of Cosine scheduler is set to 0.03. The pre-training corpus is trained for one epoch, the instruction data is trained for two epochs.
@@ 4.1 XMedBench: Multilingual Medical Knowledge Evaluation
Construction of XMedBench
For English, we use the MedQA-USMLE (Zhang et al., 2018), MedMCQA (Pal et al., 2022), and medical-related parts of MMLU (Hendrycks et al., 2020); for Chinese, we used the of MedQA-MCMLE (Zhang et al., 2018) and medical-related parts CMMLU (Li et al., 2023a); for Spanish, we used HEAD-QA (Vilares and Gómez-Rodríguez, 2019); for French, we used FrenMedMCQA (Labrak et al., 2023a); For Arabic and Hindi, which lack local evaluation, we compromised and followed Llama3’s multilingual evaluation method (et al., 2024), using Google Translate and inviting practicing physicians to proofread and finally get the translated version of MMLU. Specifically, we follow Med-PaLM2 (Singhal et al., 2023) and select six subcategories in MMLU: Clinical knowledge, Medical genetics, Anatomy, Professional medicine, College biology, and College medicine. For MedQA, we choose the 4-options version. For CMMLU, we select seven subdirectories: Anatomy, Clinical knowledge, College medicine, Genetics, Nutrition, Traditional chinese medicine, and Virology.
Settings
We adopt 3-shot evaluation and use regular matching to extract options. The specific evaluation prompts are shown in Fig. 3. For the generation strategy, we do not perform sampling and set the maximum and minimum number of generated tokens to 128 and 2. For model loading, except for the 0.5B size model which uses full precision loading, we uniformly use half precision loading. Please see the App. B.2 for details of Models.
@@
Language English Chinese French Spanish Arabic Hindi Avg.
[…]
Our Models and their Bases
Qwen-0.5B 24.43 3.78 16.94 14.16 10.88 23.68 26.02 26.29 26.35 19.17
Qwen-1.8B 26.79 31.05 40.89 44.28 35.86 22.12 27.43 23.71 8.82 28.99
Gemma-2B 30.24 32.27 37.35 25.98 28.06 25.86 32.43 20.96 25.53 28.74
Apollo-2B 38.33 42.00 52.89 46.76 36.76 38.32 41.28 31.62 31.50 39.94
Yi-6B 45.48 47.98 62.27 78.90 69.47 45.79 47.01 12.22 10.74 46.65
Apollo-6B 56.25 57.53 68.65 85.52 72.62 51.71 58.47 33.46 33.61 57.54
Gemma-7B 53.42 50.94 70.15 48.95 43.29 57.63 62.79 36.21 48.58 52.44
Apollo-7B 56.00 58.21 71.86 72.36 59.04 60.44 63.73 41.82 45.55 58.78
Table 5: Model performance comparison before and after Mix Training
@@ 5 The Application of the Lite Apollo: Proxy-Tuning for Larger Models
Preliminaries
Inspired by  Liu et al. (2024a, 2021), we introduce a lightweight model-agnostic decoding method in medical senarios. We leverage the logits from both pre and post fine-tuned small models to indirectly steer the larger base model’s adjustments, thereby eschewing the need for direct parameter fine-tuning. Let $M_{raw}$ denote the smaller pre-trained model, and $M_{tuned}$ denote its fine-tuned counterpart. We compute the logit offset as "proxy" for each token, corresponding to the anti-expert and expert roles as delineated in  Liu et al. (2021). This offset is then applied to the base model $M_{base}$ to synchronize the predictive distributions of the smaller and larger models. The modified probability distribution is given by:
@@
Settings
$M_{base}$ is designated as the subject of investigation for Qwen-7B. Apollo-1.8B and Qwen-1.8B are appointed as the $M_{tuned}$ and $M_{raw}$.
Results
As shown in the Tab. 6, the overall effect of the model improves a lot without changing the parameters after proxy-tuning. From language perspective, except English, all other languages increase, and French has the most obvious increase. Excitingly, for French and Spanish, the model after proxy-tuning performs better than both $M_{tuned}$ and $M_{base}$, indicating that new accurate knowledge is generated after proxy-tuning. We also notice a decline in English proficiency. This may be because there is a gap between the distribution of difference and the probability itself, which leads to over-strengthening of the second option and requires further exploration and optimization.